X1 Let x = V = and v2 - and let T: R2R2 be a linear transformation that maps x into xxv, + XxV2. Find a matrix A such that T(Assume that is a linear transformation. Find the standard matrix of T. T:R3-R2, T(41) = (1,3), and T(62) =(-4,6), and T(03) =Let T: R2-R2 be a linear transformation such that T(x1,x2) = (x2 + x2, 3x2 + 2x2). Find x such that T(x) = (5,18). X=Assume that is a linear transformation. Find the standard matrix of T. 51 TR2_R2, rotates points (about the origin) through -
Answer #1

-8 x = NG 22 = q= -6 27 2 T: RTR T(X) = 4 4 + x rh 8 Then the metrix Correspondand to X is Associated T(!)- 2 Mi) = V2 -7 G -w T: R² R be lineas transformat T(4,8) (26+22 32 +222) TW) = (5,78) ( 2+2 , 37,+233) 5=xtų Els 18= 34+242 (ii) را فو - ) 8 =